OpenStudy (anonymous):

Find the point (x,y) on the line y=x that is equidistant from the points (4,-1) and (7,-6)

OpenStudy (amistre64):

in other words, we need to find the line that forms thru the midpoint of the line segment created from these points and perpendicular to it; then equate it to y=x

OpenStudy (amistre64):

midpoint is add em up and divide by 2 (4,-1) (7,-6) ------ (11,-7)/2 = (11/2, -7/2) the slope between them is subtract the points and stack y/x (7,-6) -(4,-1) ------- 3,-5 ; slope = y/x = -5/3; our perp slope is flip and negate: 3/5 now we form the equation of the perp line y = (3/5)x -3(11)/5(2) - 7/2 y = (3/5)x -68/10 and when does this equate with y=x? y = (3/5)x -68/10 = y = x (3/5)x -68/10 = x -68/10 = (2/5)x -17 = x .... and by default, since y = x, when y = -17 too (-17,-17)

OpenStudy (amistre64):

we can dbl check with pythag thrm (4,-1) (-17,-17) (7,-6) 17 17 17 17 17 17 --------------------- 21,16 0 , 0 24,11 sqrt(21^2+16^2) = sqrt(24^2 + 11^2) ?? http://www.wolframalpha.com/input/?i=sqrt%2821^2%2B16^2%29+%3D+sqrt%2824^2+%2B+11^2%29

OpenStudy (anonymous):

goodlooks i get it now i had no idea what the question asked but i get it now thanks

OpenStudy (amistre64):

youre welcome :)